OpenStudy (anonymous):

what are the steps for "find the x-intercept and the y-intercept of the graph of 6x-4y=36

OpenStudy (anonymous):

To find x-ints: let y=0 and solve for x To find y-ints: let x=0 and solve for y

OpenStudy (anonymous):

x-int: 6x-4(0)=36 6x=36 x=6 y-int: 6(0)-4y=36 -4y=36 y=-9